OpenStudy (annie12m):

One hundred twenty is what percent of 200?

OpenStudy (annie12m):

@lolacole12

OpenStudy (lolacole12):

Can you fan me are you in 7th grade

OpenStudy (annie12m):

already am your fan and i am in 7th

OpenStudy (anonymous):

it is 60% since it is 200

OpenStudy (anonymous):

let me explain why

OpenStudy (anonymous):

divide it by 2

OpenStudy (annie12m):

ty @Liam2004

OpenStudy (anonymous):

and 120 becomes 60 and it then that makes the answer 60%

OpenStudy (anonymous):

thx for medal man

OpenStudy (lolacole12):

One hundred twenty is 60% of 200. 120/200 = x/100; 0.6 = x/100; x = 0.6 * 100; x = 60

OpenStudy (lolacole12):

late sorry